OpenStudy (anonymous):

YsinX+e^x=Y^2 Implicit differentiation?

OpenStudy (anonymous):

ahh now we have a question in this case you are thinking of \(y\) as a function of \(x\) although you don't know what it is (that is why it is "implicit" i.e. it is implied that \(y\) is a function of \(x\) start with the product rule \[y'\sin(x)-y\cos(x)+e^x=2yy'\] and solve for \(y'\)

OpenStudy (anonymous):

i got this part and y'=(-ycosX-e^x)/9SinX-2y but i just wanna know the whole thing cause i got a really nasty answer at the end